Circular No. 8597

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2005 R2 (VAN NESS) M. E. Van Ness, Lowell Observatory, reports his discovery of a comet

on images taken by himself with the LONEOS telescope (discovery observation tabulated below); the new object shows a moderately condensed elongated coma of diameter 20′′ with a fan-shaped tail 245′′ long in p.a. 240^{o}. Following posting on the ‘NEO Confirmation Page’, several other observers have noted the object’s cometary nature on their CCD images — including D. Kyrylenko, O. Gerashchenko, V. Lokot, and Y. Ivashchenko (Zhytomyr reg, Ukraine, Sept. 10.8 UT; 2′ tail in p.a. 250^{o}); E. Reina Lorenz (Masquefa, Spain, Sept. 11.0; 50′′ coma and 100′′ tail); C. Jacques and E. Pimentel (Belo Horizonte, Minas Gerais, Brazil, Sept. 11.25; elongated coma of diameter 9′′ × 12′′ and red mag 15.8–16.1 with a broad 50′′ fan tail in p.a. 240^{o}); and J. E. McGaha (Tucson, AZ, Sept. 11.27–11.29; six 60-s co-added images show a 12′′ × 18′′ coma elongated with a narrow, fan-shaped tail 115′′ long in in p.a. 236^{o}).

2005 UT  α_{2000}  δ_{2000}  Mag.
Sept.10.40969  2^{h}26^{m}37.^{s}95  +28^{o}36′18′′.9  17.0

The available astrometry (including prediscovery LINEAR observations on Aug. 16 and 31), the following orbital elements, and an ephemeris appear on MPEC 2005-R54.

T = 2005 Feb. 11.022 TT  ω = 3.^{o}215
e = 0.38393  Ω = 312.836 }2000.0
q = 2.12818 AU  i = 10.246
a = 3.45442 AU n^{o} = 0.153512 P = 6.42 years

COMET P/2004 FY_{140} (LINEAR) An apparently asteroidal object discovery by LINEAR (discovery ob-

servation below, from MPS 104067; orbital elements on MPO 66354) has been found by C. Hergenrother, Lunar and Planetary Laboratory, to have been consistently non-stellar on individual and co-added R-band frames taken on 2004 May 19 and 20 UT; images taken on 2005 July 5 and 6 with the Catalina 1.54-m reflector failed to show the object for 15′ along the line-of-variation (limiting mag V ∼ 22; predicted mag near V ∼ 19).

2004 UT  α_{2000}  δ_{2000}  Mag.
Mar. 27.33098  12^{h}37^{m}49.^{s}21  −6^{o}08′44′′.6  18.1
2005 September 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green